OpenStudy (anonymous):

Describe the transformation of the graph of f into the graph of g as either a horizontal or vertical stretch. \[f(x)=\sqrt{x}\] \[g(x)=3\sqrt{x}\]

OpenStudy (agent0smith):

\[\Large f(x) = a \sqrt x\] if a > 1, the graph will be stretched vertically, and 3 > 1...

OpenStudy (anonymous):

can you explain a little bit more please?

OpenStudy (anonymous):

\[f(x)=\sqrt{x}\] \[g(x)=3\sqrt{x}\]

OpenStudy (anonymous):

those equations have to be described

OpenStudy (agent0smith):

\[\Large f(x) = \sqrt x\] \[\Large f(x) = a \sqrt x\] To compare them, if a > 1, the graph will be stretched vertically, and 3 > 1, so...